# Algorithms and Complexity of Hedge Cluster Deletion Problems

### v1

@@ Abstract
A hedge graph is a graph whose edge set has been partitioned into groups called hedges. Here we consider a generalization of the well-known Cluster Deletion problem, named Hedge Cluster Deletion. The task is to compute the minimum number of hedges of a hedge graph so that their removal results in a graph that is isomorphic to a disjoint union of cliques. From the complexity point of view, unlike the classical Cluster Deletion problem, we prove that Hedge Cluster Deletion is NP-complete on non-trivial graphs which contain all graphs having unbounded size of vertex-disjoint 3-vertex-paths as subgraphs. To complement this, we show that for graphs that contain bounded size of vertex-disjoint 3-vertex-paths as subgraphs, Hedge Cluster Deletion can be solved in polynomial time. Regarding its approximability, we prove that the problem is tightly connected to the related complexity of the Min Horn Deletion problem, a well-known boolean CSP problem. Our connection shows that it is NP-hard to approximate Hedge Cluster Deletion within factor $2^{O(\log^{1-\epsilon}r)}$ for any $\epsilon>0$, where $r$ is the number of hedges in a given hedge graph.
@@ 1 Introduction
Hedge graphs represent situations where a set of edges might fail together due to some underlying dependency. In particular, a hedge graph is a graph whose edges are partitioned into groups called hedges. Motivated by the fact that edge failures might appear simultaneously, hedge graphs appear naturally in a network design problem, known as Hedge Cut. In this problem, the task is to remove a minimum number of hedges that disconnects a given hedge graph. Indeed, in a series of SODA papers there has been an increasing interest in identifying whether Hedge Cut is polynomial-time solvable or quasipolynomial-time solvable or NP-complete [ChandrasekaranX18, sodaFominGKL025, sodaFominGKLS23, sodaGhaffariKP17, JaffkeLMPS23].
@@ 1.1 Related work
As already mentioned, hedge graphs were inspired by the Hedge Cut and close related problems [ChandrasekaranX18, sodaFominGKL025, sodaFominGKLS23, sodaGhaffariKP17, JaffkeLMPS23]. Interestingly, several hedge variants of the problem were discussed. However, they mainly focus on connectivity problems in order to understand the classical complexity of Hedge Cut. We note here the Hedge Minimum Cycle problem in which the task is to detect a cycle that contains a minimum number of hedges. As the hedge function (i.e., the number of hedges covering an edge subset) is a monotone submodular function, there exists a quasipolynomial-time algorithm for the Hedge Minimum Cycle problem [sodaFominGKLS23].
In the Cluster Deletion problem we seek to delete the minimum number of edges of a given graph such that the resulting graph is a vertex-disjoint union of cliques (cluster graph). It is known that the problem is NP-hard on general graphs [CluNP04]. Settling its complexity status has attracted several researchers. If the input graph is restricted on well-known graph classes the problem remains NP-hard [BDM15, CD-cographs, GolovachHKLP20, CD-lowdegree, KonstantinidisP21], although there are graph classes where the problem is polynomial solvable [BonomoDNV15, BDM15, CD-cographs, GruttemeierK20, CD-lowdegree, KonstantinidisP21]. In Table 1, we summarize the complexity of the problem on the corresponding graph classes. Cluster Deletion has been studied extensively from the perspective of approximation algorithms. A greedy approach of selecting a maximum clique provides a $2$-approximation algorithm, though not necessarily in polynomial time [DessmarkJLLP07]. Other constant factor approximation algorithms for Cluster Deletion were based on rounding a linear programming (LP) relaxation [CGW03, PuleoM15, Veldt22a, VeldtGW18], with the current-best approximation factor of 2 achieved in [VeldtGW18]. As Cluster Deletion is an $H$-free Edge Deletion problem with $H=P_{3}$, we also mention complexity results regarding hardness of approximation for almost every fixed graph $H$ [BliznetsCKP18]. In particular, if $H$ is a 3-connected graph with at least two non-edges, or a cycle on at least four vertices or a path on at least five vertices, then $H$-free Edge Deletion does not admit poly(OPT)-approximation unless P = NP [BliznetsCKP18].
Graphs Cluster Deletion Hedge Cluster Deletion
Bipartite Poly-time (trivial)
Interval Poly-time [KonstantinidisP21]
Cograph Poly-time [CD-cographs] NP-hard (Proposition 3.2)
@@
Table 1: Complexity of Cluster Deletion and Hedge Cluster Deletion restricted on particular graph classes. Our result given in bold from Proposition 3.2 implies that in almost all non-trivial classes of graphs Hedge Cluster Deletion remains NP-hard.
@@ 1.2 Overview of our results and techniques
Here we provide a high-level description of our results. For necessary terminology and formal definitions concerning specially notions and terms used on hedge graphs, we refer to Section 2.
@@
As Hedge Cluster Deletion is a generalization of Cluster Deletion, it is natural to identify classes of graphs with the same or different complexity behavior. We begin by considering the underlying graph structure of a hedge graph. Clearly, if Cluster Deletion is NP-complete on a certain graph class, so does Hedge Cluster Deletion on the same class. Our results imply that in a wide range of non-trivial graph classes Hedge Cluster Deletion is NP-complete, whereas Cluster Deletion admits polynomial-time solution, highlighting a complexity different behavior (Proposition 3.2). More precisely, we show that if there is a subgraph that contains an unbounded number of disjoint $P_{3}$ then Hedge Cluster Deletion is NP-complete. Such graphs contain long paths or trees of constant height or almost cliques (graphs obtained from a clique by removing a single edge). Thus, in almost every interesting graph class, Hedge Cluster Deletion remains NP-complete.
@@
Here we prove that Hedge Cluster Deletion is Min Horn Deletion-complete under $A$-reductions, i.e., corresponding reductions required for the Min Horn Deletion-completeness. Intuitively, $A$-reductions preserve approximability problems up to a constant factor (or higher). For the exact meaning of related notions, we refer to Section 4. We point out that showing Min Horn Deletion-completeness, neither rules out the possibility of the existence of poly(OPT)-approximation nor shows that such an approximation exists. Despite this fact, few problems are classified as Min Horn Deletion-complete. Interestingly, the edge-deletion problem $(K_{d}-e)$-Edge Deletion was shown to be Min Horn Deletion-complete in which the goal is to delete as few edges as possible from a given graph so that the resulting graph does not contain an induced subgraph isomorphic to $K_{d}$ without an edge, for any $d\geq 5$ [BliznetsCKP18]. Subsequently, Belova and Bliznets [BelovaB22] showed that the editing variation of the latter problem still remains Min Horn Deletion-complete.
@@
Tractable results
Moreover, we reveal a connection between the subgraph of disjoint $P_{3}$ characterization and the complexity of the problem. Such an underlying subgraph characterization provides us with a complete dichotomy for Hedge Cluster Deletion with respect to NP-completeness and polynomial-time algorithms. In particular, we show that if the largest subgraph of disjoint $P_{3}$ is bounded, then Hedge Cluster Deletion is polynomial-time solvable (Theorem 3.3). For doing so, we prove that the graph obtained after removing the particular subgraph contains only an induced matching and isolated vertices. Then, by enumerating all subsolutions on the bounded-size subgraph, we are able to bound all cluster subgraphs. As a representative family of graphs that are applicable for the polynomial-time algorithm, we only mention complete bipartite graphs with bounded one of the two sizes, which admit trivial solution for Cluster Deletion.
@@
Based on the classified (in)approximability and the hardness of Hedge Cluster Deletion imposed by the structure of the underlying graph, we consider structural insights for establishing approximation guarantees or polynomial-time algorithms. In order to achieve a constant factor approximation, we restrict ourselves to hedge graphs in which no triangle is covered by more than two hedges. By our previous results (Proposition 3.2), we note that the problem still remains NP-complete even in this special class of hedge graphs. However, we now exploit an interesting property related to Hedge Cluster Deletion. Whenever a hedge is removed, we know how to compute the set of hedges that span new $P_{3}$ because of the hedge removal (Lemma 4.11). For a hedge $x$, we denote by $R(x)$ all hedges that inherit such property. We show that not only we can compute $R(x)$ in polynomial time, but there is a structural property behind these sets: (P1) for any $x\in R(y)$, we have $R(x)\subseteq R(y)$ (Lemma 4.10).
@@
In order to understand further implications that might lead to the tractability of Hedge Cluster Deletion, we consider the natural underlying structure formed by the hedges. Our goal is to exploit properties of the hedge intersection graph that are sufficient to deduce a polynomial-time algorithm for Hedge Cluster Deletion. Here we make a first step towards this direction by developing a polynomial-time algorithm for Hedge Cluster Deletion whenever the hedge intersection graph is acyclic.
Before reaching the details of our algorithm, let us mention some non-trivial classes of hedge graphs with unbounded number of hedges that admit acyclic hedge intersection graph. As a first example, consider any hedge graph with the edges in each biconnected component belonging to at most two hedges and all other edges belong to a distinct hedge. For densest graphs, consider a (subgraph of a) clique which is formed by a matching of unbounded size where each edge of the matching belongs to a distinct hedge and all other edges of the clique belong to the same hedge. Thus, there are some interesting classes of hedge graphs that admit acyclic hedge intersection graph.
The hedge intersection graph $\mathcal{F}$ of a hedge graph has its vertex set the set of hedges and two hedges in $\mathcal{F}$ are adjacent if the corresponding hedges contain a common vertex in the original hedge graph. We note that since we consider $\mathcal{F}$ to be acyclic, at most two hedges cover any triangle of the hedge graph. We also note that several properties explained earlier hold in this situation as well. In particular, sets $R(x)$ can be computed efficiently and property (P1) is fulfilled. However, here we have to be careful because between hedges in $\mathcal{F}$ there are edges caused by triangles.
It is not difficult to see that between two hedges in $\mathcal{F}$ there are two types of edges: (i) edges formed by hedges that cover only triangles or (ii) edges formed by hedges that cover only $P_{3}$. Of course, there are edges that belong to both types. In fact, for the latter type of edges we can deduce which of the two endpoints (hedge of the original hedge graph) should belong to any solution (Lemma 5.1). This is based on the fact that in any triangle one of the two hedges dominate the other, meaning that the removal of the one of the two hedges always cause a $P_{3}$.
@@ 2 Preliminaries
A clique of $G$ is a set of pairwise adjacent vertices of $G$, and a maximal clique of $G$ is a clique of $G$ that is not properly contained in any clique of $G$. An independent set of $G$ is a set of pairwise non-adjacent vertices of $G$. The complete graph on $n$ vertices is denoted by $K_{n}$ and the chordless path on three vertices is denoted by $P_{3}$.
@@
To avoid heavy notation and without affecting consistency, we consider a hedge graph $H$ as a simple graph (which corresponds to the underlying graph of H{H}) in which the edges of $E(H)$ have been partitioned into $\ell$ hedges $E_{1},\ldots,E_{\ell}$, denoted by $H=(V,\mathcal{E})$, where $\mathcal{E}(H)=\{E_{1},\ldots,E_{\ell}\}$. Thus, each edge of $H$ belongs to exactly one hedge $E_{i}$. The removal of a hedge $E_{i}$ results in a hedge graph that we denote by $H\setminus E_{i}$ which is obtained from the graph $H$ by removing all edges of $E_{i}$ and consists of the hedge set $(E_{1},\ldots,E_{\ell})\setminus E_{i}$.
@@
A natural difference that appears in the hedge variation of the problem, comes from the fact that maximal cluster subgraphs are enough to consider for the Cluster Deletion problem. However, in Hedge Cluster Deletion one has to explore all cluster hedge-subgraphs rather than the maximal ones. Given a subgraph $H^{\prime}$ with edge set $E(H^{\prime})\subseteq E(H)$, we can identify if $H^{\prime}$ is a valid solution for Hedge Cluster Deletion, i.e., whether $H^{\prime}$ being a cluster hedge-subgraph of $H$. In fact, the following algorithm implies that if the number of cluster subgraphs of $H$ is polynomial then Hedge Cluster Deletion can be solved in polynomial time: for each cluster subgraph of the underlying graph, we realize if it is indeed a cluster hedge-subgraph of the given hedge graph and keep the valid solution with a maximum number of hedges.
@@
We begin by checking if $H^{\prime}$ is a cluster graph in $O(m+n)$ time. Then we search for any non-valid hedge $E_{i}\in\mathcal{E}(H)$ with the following properties: $E_{i}\cap E(H^{\prime})\neq\emptyset$ and $E_{i}\cap(E(H)\setminus E(H^{\prime}))\neq\emptyset$. If such a hedge exists, then $H^{\prime}$ is not a hedge subgraph of $H$, since there is an edge in $E(H^{\prime})$ that does not belong to any hedge that is fully contained in $H^{\prime}$. Thus, if all hedges of $H$ are valid in $H^{\prime}$, then $H^{\prime}$ is indeed a hedge-subgraph of $H$, which means that $H^{\prime}$ is a valid solution for Hedge Cluster Deletion. ∎
Since cluster graphs are exactly the graphs that do not contain any $P_{3}$, we distinguish such paths with respect to the hedges of $H$. We say that a hedge $E_{i}$ spans an internal $P_{3}$ if there is a $P_{3}$ in $H$ having both of its edges in $E_{i}$. Otherwise, we refer to the induced $P_{3}$ of $H$ as a $P_{3}$-hedge, which means that two hedges are involved in the $P_{3}$ of $H$. By definition, we have the following.
@@
3 Hardness result and complexity dichotomy

### v2

@@ Abstract
A hedge graph is a graph whose edge set has been partitioned into groups called hedges. Here we consider a generalization of the well-known Cluster Deletion problem, named Hedge Cluster Deletion. The task is to compute the minimum number of hedges of a hedge graph so that their removal results in a graph that is isomorphic to a disjoint union of cliques. From the complexity point of view, unlike the classical Cluster Deletion problem, we prove that Hedge Cluster Deletion is NP-complete on non-trivial graphs which contain all graphs having a large size of vertex-disjoint 3-vertex-paths as subgraphs. To complement this, we show that for graphs that contain bounded size of vertex-disjoint 3-vertex-paths as subgraphs, Hedge Cluster Deletion can be solved in polynomial time.
Regarding its approximability, we prove that the problem is tightly connected to the related complexity of the Min Horn Deletion problem, a well-known boolean CSP problem. Our connection shows that it is NP-hard to approximate Hedge Cluster Deletion within factor $2^{O(\log^{1-\epsilon}r)}$ for any $\epsilon>0$, where $r$ is the number of hedges in a given hedge graph. We also address the parameterized complexity of the problem. While Hedge Cluster Deletion is fixed-parameter tractable with respect to the solution size (i.e., the number of removal hedges), we prove that it does not admit a polynomial kernel, unless NP $\subseteq$ coNP/poly.
@@ 1 Introduction
Edge deletion problems are motivated by applications in denoising data derived from imprecise experimental measurements and, therefore, have been extensively studied from classical, approximation, and parameterized complexity. The objective is usually concerned with the property of being $H$-free graph for some fixed graph $H$ and almost a complete classification with respect to $H$ is known [KratschW13, CaiC15, BelovaB22, BliznetsCKP18, AravindSS17]. However, little is known when generalizing the concept from edge deletion problems to set deletion problems. Considering groups of edges as potential sets that can be removed is captured by the notion of hedge graphs.
Hedge graphs represent situations where a set of edges might fail together due to some underlying dependency. In particular, a hedge graph is a graph whose edges are partitioned into groups called hedges. Motivated by the fact that edge failures might appear simultaneously, hedge graphs appear naturally in a network design problem, known as Hedge Cut. In this problem, the task is to remove a minimum number of hedges that disconnects a given hedge graph. Indeed, in a series of recent papers there has been an increasing interest in identifying whether Hedge Cut is polynomial-time solvable or quasipolynomial-time solvable or NP-complete [ChandrasekaranX18, sodaFominGKL025, sodaFominGKLS23, sodaGhaffariKP17, JaffkeLMPS23].
@@ 1.1 Related work
As already mentioned, hedge graphs were inspired by the Hedge Cut and closely related problems [ChandrasekaranX18, sodaFominGKL025, sodaFominGKLS23, sodaGhaffariKP17, JaffkeLMPS23]. Interestingly, several hedge variants of the problem were discussed. Their main focus is on connectivity problems in order to understand the classical complexity of Hedge Cut. A quasipolynomial-time algorithm was proposed for detecting a cycle of minimum submodular cost on multigraphs [sodaGhaffariKP17, FominGKLS24]. As the partitioning function into hedges is submodular (i.e., the number of hedges covering an edge subset), the same algorithm can be applied for the Hedge Minimum Cycle problem in which the task is to detect a cycle that contains a minimum number of hedges [sodaGhaffariKP17, FominGKLS24]. However, the hedge cut function is not necessarily submodular [sodaGhaffariKP17]. A recent approach that overcomes such difficulty related to hedge cut problems uses alternative measures of connectivity and proposes a polymatroid perspective of hedge graphs [abs-2510-25043].
In the Cluster Deletion problem we seek to delete a minimum number of edges of a given graph such that the resulting graph is a vertex-disjoint union of cliques (cluster graph). It is known that the problem is NP-hard on general graphs [CluNP04]. Settling its complexity status has attracted several researchers. If the input graph is restricted on well-known graph classes the problem remains NP-hard [BDM15, CD-cographs, GolovachHKLP20, CD-lowdegree, KonstantinidisP21], although there are graph classes where the problem is polynomial solvable [BonomoDNV15, BDM15, CD-cographs, GruttemeierK20, CD-lowdegree, KonstantinidisP21]. In Section 1.1, we summarize the complexity of the problem on the corresponding graph classes. Cluster Deletion has been studied extensively from the perspective of approximation algorithms. A greedy approach of selecting a maximum clique provides a $2$-approximation algorithm, though not necessarily in polynomial time [DessmarkJLLP07]. Other constant factor approximation algorithms for Cluster Deletion were based on rounding a linear programming (LP) relaxation [CGW03, PuleoM15, Veldt22a, VeldtGW18], with the current-best approximation factor of 2 achieved in [VeldtGW18].
[…]
Restriction (graph classes) Cluster Deletion Hedge Cluster Deletion
NP-hard or poly-time Bipartite Poly-time (trivial)
Interval Poly-time [KonstantinidisP21] NP-hard
Cograph Poly-time [CD-cographs] (Theorem 3.2)
@@
Special bipartite1 Poly-time (trivial) Poly-time
Special split2 Poly-time [BDM15] (Theorem 3.3)
Acyclic hedge intersection Poly-time (trivial) Poly-time (Theorem 6.7)
Approxim. guarantees No restriction Constant factor [VeldtGW18] Min Horn Deletion-complete (Theorem 4.7)
Bi-hedge graphs        – Constant factor (Theorem 5.7)
Parameter. complexity Parameterized by the solution size Polynomial kernel [CaoC12, GuillemotHPP13] No polynomial kernel (Theorem 4.12)
[…]
@@ 1.2 Overview of our results and techniques
Here we provide a high-level description of our results. Section 1.1 summarizes our main results. For necessary terminology and formal definitions concerning specially notions and terms used on hedge graphs, we refer to Section 2. However, in the following overview it is enough to recall that the edges of a hedge graph are partitioned into hedges and the underlying graph corresponds to the simple graph obtained from a hedge graph by ignoring the hedges.
@@
As Hedge Cluster Deletion is a generalization of Cluster Deletion, it is natural to identify classes of graphs with the same or different complexity behavior. We begin by considering the underlying graph structure of a hedge graph. Clearly, if Cluster Deletion is NP-complete on a certain graph class, so is Hedge Cluster Deletion on the same class. Our results imply that in a wide range of non-trivial graph classes Hedge Cluster Deletion is NP-complete, whereas Cluster Deletion admits polynomial-time solution, highlighting a complexity different behavior (Theorem 3.2). More precisely, we show that if there is a subgraph that contains a large enough number of vertex-disjoint $P_{3}$’s then Hedge Cluster Deletion is NP-complete. Such graphs contain long paths or trees of constant height or almost cliques (graphs obtained from a clique by removing a single edge). Thus, in almost every interesting graph class, Hedge Cluster Deletion remains NP-complete.
Tractable results
Moreover, we reveal a connection between the subgraph of disjoint $P_{3}$ characterization and the complexity of the problem. In particular, we show that if the largest subgraph of vertex-disjoint $P_{3}$’s is bounded, then Hedge Cluster Deletion is polynomial-time solvable (Theorem 3.3). For doing so, we prove that the graph obtained after removing the particular subgraph contains only an induced matching and isolated vertices. Then, by enumerating all subsolutions on the bounded-size subgraph, we are able to bound all cluster subgraphs. As a representative family of graphs that are applicable for the polynomial-time algorithm, we only mention complete bipartite graphs with bounded one of the two sizes, which admit trivial solution for Cluster Deletion.
@@
Here we prove that Hedge Cluster Deletion is Min Horn Deletion-complete under $A$-reductions, i.e., corresponding reductions required for the Min Horn Deletion-completeness. Intuitively, $A$-reductions preserve approximability problems up to a constant factor (or higher). For the exact meaning of related notions, we refer to Section 4.1. We point out that showing Min Horn Deletion-completeness, neither rules out the possibility of the existence of poly(OPT)-approximation nor shows that such an approximation exists. Despite this fact, few problems are classified as Min Horn Deletion-complete. Interestingly, the edge-deletion problem $(K_{d}-e)$-Edge Deletion was shown to be Min Horn Deletion-complete in which the goal is to delete as few edges as possible from a given graph so that the resulting graph does not contain an induced subgraph isomorphic to $K_{d}$ without an edge, for any $d\geq 5$ [BliznetsCKP18]. Subsequently, Belova and Bliznets [BelovaB22] showed that the editing variation of the latter problem still remains Min Horn Deletion-complete.
@@
Incompressibility
Interesting connections arise with respect to parameterized complexity. Recall that Hedge Cluster Deletion is easily shown to be FPT parameterized by the solution size with running time $2^{k}n^{O(1)}$. In fact, the running time is essentially the best one can hope for because Cluster Deletion has no subexponential-time algorithm [CD-lowdegree]. However, polynomial kernels for Cluster Deletion do exist [CaoC12, GuillemotHPP13]. Therefore, it is natural to ask whether Hedge Cluster Deletion admits a polynomial kernel. We answer negative to this question under reasonable hierarchy assumptions, which highlights another complexity difference between the two problems.
More general, there seems to be a connection between the existence of polynomial kernel and poly(OPT)-approximation for a problem, because usually one can find poly(OPT)-approximation having polynomial kernel, and vice versa. Although this fact is not formal and there are problems that act differently [GiannopoulouLSS16], it is quite possible that such connection exists for edge deletion problems: up to now, the existence of poly(OPT)-approximation algorithms are similar to results for kernelization complexity [BliznetsCKP18, BelovaB22, CaiC15]. We establish the incompressibility of Hedge Cluster Deletion by ppt-reduction from an incompressible constraint satisfiability problem (Theorem 4.12). This specific problem describes propagational behaviors on Boolean variables and clauses after certain events occur, which encompasses the situation in hedge graphs after each hedge removal. We do so, by representing Boolean formulas by three-vertex components that ensure the consistency of $P_{3}$-freeness in a similar way to the MinOnes()($\mathcal{F}$) problem. Interestingly, we also show a connection between the MinOnes()($\mathcal{F}$) problem and the incompressible constraint satisfiability problem via the Hedge Cluster Deletion problem (Theorem 7.1).
@@
Based on the classified (in)approximability and the hardness of Hedge Cluster Deletion imposed by the structure of the underlying graph, we consider structural insights for establishing approximation guarantees or polynomial-time algorithms. In order to achieve a constant factor approximation, we restrict ourselves to hedge graphs in which no triangle is covered by more than two hedges. By our previous results (Theorem 3.2), we note that the problem still remains NP-complete even in this special class of hedge graphs. However, we now exploit an interesting property related to Hedge Cluster Deletion. Whenever a hedge is removed, we know how to compute the set of hedges that span new $P_{3}$ because of the hedge removal (Lemma 5.2). For a hedge $x$, we denote by $R(x)$ all hedges that inherit such property. We show that not only we can compute $R(x)$ in polynomial time, but there is a structural property behind these sets: (P1) for any $y\in R(x)$, we have $R(y)\subseteq R(x)$ (Lemma 5.1).
@@
In order to understand further implications that might lead to the tractability of Hedge Cluster Deletion, we consider the natural underlying structure formed by the hedges. Our goal is to exploit properties of the hedge intersection graph that are sufficient to deduce a polynomial-time algorithm for Hedge Cluster Deletion. Here we take a first step towards this direction by developing a polynomial-time algorithm for Hedge Cluster Deletion whenever the hedge intersection graph is acyclic.
Before reaching the details of our algorithm, let us mention some non-trivial classes of hedge graphs with unbounded number of hedges that admit acyclic hedge intersection graph. As a first example, consider any hedge graph with the edges in each biconnected component belonging to at most two hedges and all other edges belong to a distinct hedge. For dense graphs, consider a (subgraph of a) clique which is formed by a matching of unbounded size where each edge of the matching belongs to a distinct hedge and all other edges of the clique belong to the same hedge. Thus, there are some interesting classes of hedge graphs that admit an acyclic hedge intersection graph.
The hedge intersection graph $\mathcal{F}$ of a hedge graph is defined as follows: its vertex set is the set of hedges and two hedges in $\mathcal{F}$ are adjacent if the corresponding hedges contain a common vertex in the original hedge graph. We note that since we consider $\mathcal{F}$ to be acyclic, at most two hedges cover any triangle of the hedge graph. We also note that several properties explained earlier hold in this situation as well. In particular, the sets $R(x)$ can be computed efficiently and property (P1) is fulfilled. However, here we have to be careful because between hedges in $\mathcal{F}$ there are edges caused by triangles of the hedge graph.
It is not difficult to see that between two hedges in $\mathcal{F}$ there are two types of edges: (i) edges formed by hedges that cover only triangles or (ii) edges formed by hedges that cover only $P_{3}$. Of course, there are edges that belong to both types. In fact, for the latter type of edges we can deduce which of the two endpoints (hedge of the original hedge graph) should belong to any solution (Lemma 6.1). This is based on the fact that in any triangle one of the two hedges dominates the other, meaning that the removal of one of the two hedges always cause a $P_{3}$.
@@ 2 Preliminaries
A clique of $G$ is a set of pairwise adjacent vertices of $G$, and a maximal clique of $G$ is a clique of $G$ that is not properly contained in any clique of $G$. An independent set of $G$ is a set of pairwise non-adjacent vertices of $G$. The complete graph on $n$ vertices is denoted by $K_{n}$ and the chordless path on three vertices is denoted by $P_{3}$. If the vertices $a,b,c$ induce a $P_{3}$ with $\{a,c\}\notin E(G)$, we write $P_{3}=(a,b,c)$. Since we mainly deal with $P_{3}$ and $K_{3}$, we also write $K_{3}=(a,b,c)$ to denote that the vertices $a,b,c$ induce a $K_{3}$.
@@
Kernelization lower bounds
Here we include definitions from the parameterized complexity theory and kernelization; for further details we refer to [CyganFKLMPPS15, DowneyF13]. A problem with input size $n$ and parameter $k$ is fixed parameter tractable (FPT), if it can be solved in time $f(k)\cdot n^{O(1)}$ for some computable function $f$. Respectively, the complexity class FPT is composed by all fixed parameter tractable problems. A kernelization for a parameterized decision problem $Q$ is a polynomial-time preprocessing algorithm that reduces any parameterized instance $(x,k)$ into an instance $(x^{\prime},k^{\prime})$ whose size is bounded by a function $g(k)$ and which has the same answer for $Q$. The function $g$ defines the size of the kernel and the kernel has polynomial size if $g$ is polynomial on $k$.
A polynomial compression of a parameterized problem $P$ into a (nonparameterized) problem $Q$ is a polynomial algorithm that takes as an input an instance $(x,k)$ of $P$ and returns an instance $x^{\prime}$ of $Q$ such that (i) $(x,k)$ is a yes-instance of $P$ if and only if $x^{\prime}$ is a yes-instance of $Q$ and (ii) the size of $x^{\prime}$ is bounded by $p(k)$ for a polynomial $p$. Clearly, the existence of a polynomial kernel implies that the problem admit a polynomial compression but not the other way around.
It is well-known that every decidable parameterized problem is FPT if and only if it admits a kernel, but it is unlikely that every problem in FPT has a polynomial kernel or polynomial compression. In particular, the standard composition and cross-composition techniques [BodlaenderDFH09, BodlaenderJK14] allow to show that certain problems have no polynomial compressions, unless NP $\subseteq$ coNP/poly. We will apply the following reduction, known as polynomial parameter transformation (ppt-reduction in short), which plays a central role in demonstrating incompressibility (and, hence, no polynomial kernel) [BodlaenderJK14]. A ppt-reduction from a parameterized problem $P$ to another parameterized problem $P^{\prime}$ is an algorithm that, for input $(I,k)\in P$, takes time polynomial in $|I|+k$ and outputs an instance $(I^{\prime},k^{\prime})\in P^{\prime}$, such that
  • (i)
    $(I,k)$ is a yes-instance of $P$ if and only if $(I^{\prime},k^{\prime})$ is a yes-instance of $P^{\prime}$,
  • (ii)
    parameter $k^{\prime}$ is bounded by a polynomial of $k$.
It is known that if there is a ppt-reduction from a parameterized problem $P$ to another parameterized problem $P^{\prime}$, then $P^{\prime}$ admits no polynomial compression (hence no polynomial kernel) whenever $P$ admits no polynomial compression [BodlaenderJK14].
@@
We consider a hedge graph $H$ as a simple graph (which corresponds to the underlying graph of H{H}) in which the edges of $E(H)$ have been partitioned into $\ell$ hedges $E_{1},\ldots,E_{\ell}$, denoted by $H=(V,\mathcal{E})$, where $\mathcal{E}(H)=\{E_{1},\ldots,E_{\ell}\}$. Thus, each edge of $H$ belongs to exactly one hedge $E_{i}$. The removal of a hedge $E_{i}$ results in a hedge graph that we denote by $H\setminus E_{i}$ which is obtained from the graph $H$ by removing all edges of $E_{i}$ and consists of the hedge set $(E_{1},\ldots,E_{\ell})\setminus E_{i}$.
@@
A natural difference that appears in the hedge variation of the problem, comes from the fact that maximal cluster subgraphs are enough to consider for the Cluster Deletion problem. However, in Hedge Cluster Deletion one has to explore all cluster hedge-subgraphs rather than the maximal ones. Given a subgraph $H^{\prime}$ with edge set $E(H^{\prime})\subseteq E(H)$, we can identify if $H^{\prime}$ is a valid solution for Hedge Cluster Deletion, i.e., whether $H^{\prime}$ is a cluster hedge-subgraph of $H$. In fact, the following algorithm implies that if the number of cluster subgraphs of $H$ is polynomial then Hedge Cluster Deletion can be solved in polynomial time: for each cluster subgraph of the underlying graph, we realize if it is indeed a cluster hedge-subgraph of the given hedge graph and keep the valid solution with a maximum number of hedges.
@@
We begin by checking if $H^{\prime}$ is a cluster graph in $O(m+n)$ time. Then we search for any non-valid hedge $E_{i}\in\mathcal{E}(H)$ with the following properties: $E_{i}\cap E(H^{\prime})\neq\emptyset$ and $E_{i}\cap(E(H)\setminus E(H^{\prime}))\neq\emptyset$. If such a hedge exists, then $H^{\prime}$ is not a hedge subgraph of $H$, since there is an edge in $E(H^{\prime})$ that does not belong to any hedge that is fully contained in $H^{\prime}$. Thus, if all hedges of $H$ are valid in $H^{\prime}$, then $H^{\prime}$ is indeed a hedge-subgraph of $H$, which means that $H^{\prime}$ is a cluster hedge-subgraph of $H$. ∎
Since cluster graphs are exactly the graphs that do not contain any $P_{3}$, we distinguish such paths with respect to the hedges of $H$. We say that a hedge $E_{i}$ spans an internal $P_{3}$ if there is a $P_{3}$ in $H$ having both of its edges in $E_{i}$. Otherwise, we refer to the induced $P_{3}$ of $H$ as a simple $P_{3}$, which means that two hedges are involved in the $P_{3}$ of $H$. By definition, we have the following.
@@
3 Hardness result implied by the underlying structure